x - 25y = 30 & y = -34x + 350
Answer by roseo(33)

Because one equation is already solved for y use that for substitution in the other equation.
x-25y=30
y=-34x + 350
x-25(-34x+350)=30
x+850x-8750=30
851x-8750=30
851x=8780
x=8780/851
Then substitute this value into either equation to find y. The second one looks the best.
y= -34(8780/851)+350
y= -3356/4255
y=-.787
x=8780/851
x = 10.317
When you gave the original problem the problem became much easier and the two equations that you should have written were
2A + 3B = 34 and
25A + 30 B = 350 Where A = the # of A bikes and B = the # of B bikes
then multiply the first by -10 so that you have
-20A - 30B = -340 and add that to the second
25A +30B = 350 and you have
5A = 10 and then
A = 2 and then 2(2)+3B = 34
so 4+3B = 34
3B = 30 and B = 10
(10.317,-.787)